Session: Sep 1999 (session 3)
Wavelength: 30 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Thu Sep 9 10:26:40 METDST 1999) Failure
After the failure of our old formatter, the new MkIV upgrade is not yet fully under FS control. Efforts continue to get it working and join the present observing session.

Effelsberg (A. Kraus at Thu Sep 9 11:47:01 METDST 1999) Minor failures
Data lost from UT13.00 to UT13.16 due to computer problems.

Onsala (A. Polatidis at Thu Sep 9 11:52:11 METDST 1999) Success

Westerbork (Tony Foley at Fri Sep 10 18:02:40 METDST 1999) Success
12 telescopes in tied array. SEFD about 140Jy Array re-phased about 15:04 UT

Jodrell (Alastair Gunn at Wed Sep 15 12:12:49 METDST 1999) Success
Rapidly changing RFI spikes throughout band. No known problems.

Calibration (Jean\-Francois Desmurs at Sun Nov 7 15:18:23 MET 1999) Minor failures
The ANTAB formatted file for this experiment can be found on VLBEER in the sep99 directory under the name: .antab. Read the comments in the file to get further information. (Rem: Jb no Tsys values).